Omar has decided to purchase an $11,000 car. He plans on putting 20% down toward the purchase, and financing the rest at 4.8% interest rate for 3 years. find his monthly payment.
Answer by Theo(13342)

the car is 11000.

he puts 20% down which is equal to 2200.

the balance to be paid off on installments is 8800.

the interest rate is 4.8% per year compounded monthly.

use an annuity calculator to find the required payment at the end of each month.

one such calculator can be found at https://arachnoid.com/finance/

results from this calculator are shown below:

$$$

your inputs to this calculator would be:

present value = 8800
future value = 0
number of periods = 3 * 12 = 36
payment = 0 or left blank
rate percent = 4.8 / 12 = .4
payment is at end of time period

then click on PMT and you'll get the answer.